### v3

# Bridging Cache-Friendliness and Concurrency: A Locality-Optimized In-Memory B-Skiplist

Abstract.
Skiplists are widely used for in-memory indexing in many key-value stores, such as RocksDB and LevelDB, due to their ease of implementation and simple concurrency control mechanisms. However, traditional skiplists suffer from poor cache locality, as they store only a single element per node, leaving performance on the table. Minimizing last-level cache misses is key to maximizing in-memory index performance, making high cache locality essential.
In this paper, we present a practical concurrent B-skiplist that enhances cache locality and performance while preserving the simplicity of traditional skiplist structures and concurrency control schemes. Our key contributions include a top-down, single-pass insertion algorithm for B-skiplists and a corresponding simple and efficient top-down concurrency control scheme.
On $128$ threads, the proposed concurrent B-skiplist achieves between $2\times$$9\times$ higher throughput compared to state-of-the-art concurrent skiplist implementations, including Facebook’s concurrent skiplist from Folly and the Java ConcurrentSkipListMap. Furthermore, we find that the B-skiplist achieves competitive ($0.9\times$$1.7\times$) throughput on point workloads compared to state-of-the-art cache-optimized tree-based indices (e.g., Masstree). For a more complete picture of the performance, we also measure the latency of skiplist- and tree-based indices and find that the B-skiplist achieves between $3.5\times$$103\times$ lower 99% latency compared to other concurrent skiplists and between $0.85\times$$64\times$ lower 99% latency compared to tree-based indices on point workloads with inserts.
B-skiplist, blocked skiplist, concurrency
journalyear: 2025copyright: rightsretainedconference: 54th International Conference on Parallel Processing; September 08–11, 2025; San Diego, CA, USAbooktitle: 54th International Conference on Parallel Processing (ICPP ’25), September 08–11, 2025, San Diego, CA, USAdoi: 10.1145/3754598.3754655isbn: 979-8-4007-2074-1/25/09
1. Introduction
The skiplist (Pugh, 1990b) has become a widely used in-memory index (i.e., the memtable) in many popular databases, including HBase (HBase, [n. d.]), RocksDB (RocksDB, [n. d.]), and LevelDB (LevelDB, [n. d.]). Additionally, Java features a skiplist as its primary concurrent set and map implementation (Java, [n. d.]).
The main reason for the choice of skiplists (Pugh, 1990b) over trees (e.g., the B-tree (Bayer and McCreight, 1972)) is because skiplists enable simple structural modification operations. In contrast to tree-based indices, skiplists do not require complex rebalancing operations because elements are randomly (using coin tosses) assigned a height upfront. As a result, skiplists support simple and effective (both lock-based and lock-free) concurrency control (CC) schemes (Herlihy et al., 2006; Pugh, 1990a; Herlihy et al., 2007; Fomitchev and Ruppert, 2004; Fraser, 2004).
YCSB Workload Skiplist B-tree B-skiplist SL/BSL BT/BSL
(This paper)
Load + C 4.9E9 2.1E9 1.5E9 3.2 1.4
Load + E 1.1E10 2.3E9 2.0E9 5.6 1.2
Table 1. LLC load misses [^2] Section 5 contains all details about the experimental setup. of Facebook’s folly skiplist (SL) (fol, [n. d.]), a concurrent B-tree (BT) (Xu et al., 2023) and the concurrent B-skiplist (this paper, BSL) during the YCSB (Cooper et al., 2010) load and run phase. The load phase has 100% inserts, workload C has 100% finds, and E has 95% range queries/5% inserts.
Locality issues in skiplists. Unfortunately, traditional skiplists exhibit poor spatial locality because they store a single element per node. In contrast, cache-friendly indexes such as B-trees (Bayer and McCreight, 1972) store multiple elements per node, reducing the height of the structure and therefore the number of memory fetches during top-to-bottom traversals. Furthermore, storing multiple elements per node further reduces cache misses during horizontal traversals.
Table 2 shows that skiplists incur significantly more cache misses than B-trees, leaving performance on the table. Concretely, on the tested workloads, a state-of-the-art skiplist (fol, [n. d.]) incurs $2.4-4.8\times$ more cache misses than a comparable B-tree (Xu et al., 2023) on both point and range workloads. This discrepancy in cache misses translates into actual performance: the B-tree achieves between $2\times$-$8\times$ higher throughput than the skiplist on these workloads.
In this paper, we introduce a concurrent B-skiplist that improves the cache locality (and therefore the performance) of the traditional skiplist without giving up on the skiplist’s simple and effective CC schemes. Furthermore, due to its simplified CC scheme, the B-skiplist achieves lower worst-case latency than B-trees.
Skiplist structure. At a high level, skiplists (Pugh, 1990b) are randomized self-balancing data structures that support fast search, insertion, and deletion in $O(\log n)$ time with high probability [^3] An event $E_{n}$ on a problem of size $n$ occurs with high probabilty if $Pr[E_{n}]\geq 1-1/n^{c}$ for some constant $c$. (w.h.p.) by layering multiple linked lists. The bottommost level is a standard sorted linked list, while higher levels serve as “express lanes,” allowing searches to skip over multiple elements at a time.
The skiplist is parameterized by a promotion probability $p$, which determines the likelihood of elements appearing in higher levels of linked lists. Upon insertion, an element is randomly (using coin tosses with probability of heads $p$) assigned a height equal to the number of successive coin tosses until heads for that particular element. The height of an element is the number of linked-list levels that element appears in, starting from the lowest level.
Theoretical steps towards a cache-optimized skiplist. We will use the I/O model (or external-memory model) (Aggarwal and Vitter, 1988), which measures how well an algorithm or data structure takes advantage of spatial locality by measuring cache-line transfers. The model is parameterized by a cache-line size $Z$. Transferring $Z$ contiguous elements in one cache line has unit cost in the model.
Theoretically, given some node size $B=\Theta(Z)$, the straightforward way to improve the cache-friendliness of skiplists is to promote elements with probability $1/B$ rather than $1/2$ as in traditional skiplists. Indeed, Golovin (Golovin, 2010) proposed this method with the B-skiplist, a cache-optimized skiplist that matches the B-tree’s bounds in expectation. For each level in this theoretical B-skiplist, consecutive unpromoted elements are stored in the same node. However, this initial paper on B-skiplists stops short of providing theoretical guarantees w.h.p., parallelization, and implementation, leaving a massive gap in making B-skiplists practical.
Challenges to blocking skiplists. In practice, addressing locality issues in skiplists via blocking (i.e., storing multiple elements per node) raises challenges due to the randomized variable size of nodes. Theoretically, each node in a B-skiplist has $\Theta(B)$ elements in expectation, but there exist nodes with as many as $\Theta(B\log n)$ elements w.h.p. (Bender et al., 2016, 2017). Since nodes can be large, finds and inserts, which require scanning and potentially shifting a linear number of elements in a node, can cost as much as $O(\log n)$ cache-line transfers in the I/O model, matching a regular skiplist’s randomized bounds.
Bounding the element moves in a B-skiplist. To mitigate this issue in practice, we enforce fixed-size physical nodes in the B-skiplist to bound the maximum number of element moves during insertions. B-skiplist nodes are allowed to grow to arbitrary sizes according to the results of randomized promotions. However, if a logical B-skiplist node contains $k>B$ elements, we physically store it as $\lceil{k/B}\rceil$ nodes at the same level connected by pointers.
The design choice of fixed-size nodes is subtle but is key to fast inserts in practice because it limits the worst-case number of cache-line writes to $\Theta(1)$ per level. In contrast, if physical nodes are allowed to grow arbitrarily, the maximum number of cache-line writes in a B-skiplist node is $O(\log n)$ w.h.p. To be clear, this choice does not affect the overall insertion bound in the I/O model as each insert still requires $O(\log n)$ cache-line reads w.h.p. to determine the correct position. However, as we shall see in the empirical evaluation, fixed-size nodes enable the B-skiplist to support fast insertions and low variance in the latency of operations.
Prior steps towards concurrent blocked skiplists. Furthermore, any candidate for the in-memory index in databases, including the proposed B-skiplist, must be concurrent to take full advantage of parallel resources in today’s multicore machines.
On the practical side, there have been several steps towards improving locality in concurrent skiplists (Xie et al., 2017; Na et al., 2023), but these often involve periodic rebuilding of the upper levels, increasing the worst-case latency of individual operations. For example, CSSL (Sprenger et al., 2016) and PI (Xie et al., 2017) periodically rebuild the upper levels of the index, blocking reads and writes during the restructuring process.
Other blocked skiplists vary the component data structures and CC schemes at different levels, giving up on the simplicity of the original concurrent skiplists. For example, ESL (Na et al., 2023) first inserts elements only into the bottom level and later updates the upper index levels asynchronously with background threads. The ESL is composed of two levels with distinct index structures and CC mechanisms. Similarly, S3 (Zhang et al., 2019), another cache-sensitive skiplist, employs a similar strategy of a two-level index with different CC schemes. Furthermore, it adaptively chooses “guard entries” with a neural model, giving up on the randomization of skiplists and therefore the skiplist’s probabilistic theoretical guarantees.
Designing a concurrent B-skiplist. In this work, we focus on designing a simple and effective CC mechanism for B-skiplists without modifying its high-level structure. Specifically, we propose a CC mechanism based on fine-grained locking to achieve both high throughput and low worst-case latency. The ideal CC scheme for B-skiplists inherits the simplicity and performance of skiplist-based CC schemes (Pugh, 1990a; Herlihy et al., 2006; Fomitchev and Ruppert, 2004; Herlihy et al., 2007), so we use them as a starting point. However, they do not handle node splits and merges.
At a high level, skiplist insertion algorithms (and their corresponding CC mechanisms) make two traversals through the index if an element is promoted: one “top-down” read-only phase to determine the location(s) that an element should be inserted, and then a corresponding “bottom-up” write-only phase that links in the new skiplist nodes (Pugh, 1990a; Herlihy et al., 2006; Fomitchev and Ruppert, 2004; Herlihy et al., 2007; Platz et al., 2019). In both the top-down and bottom-up phase, a CC scheme for skiplists may either 1) hold a constant number of locks via hand-over-hand locking (Pugh, 1990a) or 2) hold locks on all levels that an element is promoted to (Herlihy et al., 2006; Fomitchev and Ruppert, 2004; Herlihy et al., 2007; Platz et al., 2019).
To reduce the number of traversals in both sequential and concurrent insertions, we introduce a top-down insertion [^4] We focus on the case of insertions due to space constraints, but deletions are symmetric. algorithm and that exploits inherent skiplist properties to traverse the B-skiplist only once (top-to-bottom and left-to-right). In contrast, existing skiplist insertion algorithms make two traversals if an element is promoted. As we shall see, the design of the insertion algorithm has major implications for the corresponding CC mechanism.
We build upon this top-down insertion algorithm to develop a simple yet efficient top-down CC scheme based on reader-writer locks (Courtois et al., 1971), a common synchronization primitive for in-memory indexes. The top-down CC scheme minimizes overheads by 1) avoiding multiple retires from root-to-leaf, and 2) minimizing both the number of exclusive locks held at one time and the duration that the locks are held. Using the top-down CC scheme, a thread only needs to lock nodes on at most two layers at once, and locks only a constant number of nodes at once.
Contributions. The contributions of the paper are as follows:
  • The design of the B-skiplist with fixed-size nodes to improve cache locality and mitigate worst-case behavior.
  • A novel top-down concurrency control mechanism for B-skiplists built on a corresponding insertion algorithm for B-skiplists that completes insertions in one pass from top-to-bottom.
  • A C++ implementation of the concurrent B-skiplist.
  • An empirical evaluation of the concurrent B-skiplist compared to several in-memory skiplist-based and B-tree-based indexes that demonstrates that the B-skiplist achieves between $2\times$$9\times$ higher throughput on YCSB workloads (Cooper et al., 2010) compared to state-of-the-art concurrent skiplists.
Load A B C E 11551010Workload Normalized throughput NHSFolly SLJava SLB-skiplist
Figure 1. Normalized throughput (ops/s) of skiplist-based indices relative to the No Hot Spot Skip List(NHS)  (Crain et al., 2013).
Results summary. Figure 1 demonstrates that the B-skiplist achieves between $2\times$$9\times$ higher throughput on workloads from the popular Yahoo! Cloud Serving Benchmark (YCSB) (Cooper et al., 2010) compared to state-of-the-art concurrent skiplists including Facebook’s concurrent skiplist from the Folly library (fol, [n. d.]), the Java ConcurrentSkipListMap (Java, [n. d.]), and the No Hot Spot Skiplist (NHS) (Crain et al., 2013). The YCSB workloads include a mix of point operations (finds/inserts) and range operations.
Furthermore, we also evaluate two tree-based indices: a state-of-the-art concurrent B-tree (Xu et al., 2023) and Masstree (Mao et al., 2012), a popular cache-optimized in-memory index. Compared to tree-based indices, we find that the B-skiplist achieves competitive throughput (between $0.9\times$$1.7\times$) on point workloads, and between $0.7\times$$7.5\times$ throughput on range workloads.
In addition to throughput, we evaluate all data structures on latency as well for a more complete picture of data-structure performance. As we shall detail in Section 5, the B-skiplist achieves between $3.5\times$$103\times$ lower 99% latency compared to existing concurrent skiplists.
2. Preliminaries
This section will give background on the structure and operations of skiplists and B-skiplists necessary to understand later sections. For the operations, it will focus on inserts for simplicity, but deletes are symmetric. Furthermore, it will review the reader-writer synchronization primitive, which is the core functionality that the proposed top-down single-pass concurrency control scheme for B-skiplists is based on.
Operations. A key-value dictionary data type stores pairs of keys and values (k, v). Their main operations are as follows:
  • find(k): return the associated value v.
  • insert(k, v): add (k, v) to the data structure.
  • range(k, f, length): apply the function f to the length key-value pairs with the smallest keys that are at least k.
We consider these operations since they comprise the popular YCSB (Cooper et al., 2010) workloads we use to perform the evaluation in Section 5.
2.1. Skiplist structure and operations
Figure 2. Example of a skiplist.
Structure. The skiplist (Pugh, 1990b) is a data structure that stores a hierarchy of levels of linked lists. Each linked list is sorted by the keys of the elements in that linked list, with sentinels at the beginning and end for -$\infty$ and $\infty$. The bottommost level (level 0) contains all of the elements in the data structure. The list size decreases by a constant factor in expectation at each successive level as we move up the hierarchy. Each linked list at some level $\ell>0$ contains a subset of the elements in the linked list below it (level $\ell-1$). This property is called inclusion invariant, where every element present at level $\ell$ must also be present at levels $0,\ldots,\ell-1$.
Figure 2 illustrates the skiplist’s pointer structure. Just like in trees, we will refer to the nodes at the lowest level of the skiplist ($\ell=0$) as leaf nodes and those at higher levels ($\ell>0$) as internal nodes. All nodes have a next pointer to a successor node in the same level, and all internal nodes also have a down pointer to the node with the same key in the level below. The root node is the leftmost node at the highest level.
An element that appears at some maximum level $\ell>0$ is said to be promoted to that level. Promotions are determined upfront at the start of an insertion via randomization with a series of coin flips. Notably, the level that an element is promoted to in a skiplist is unrelated to the current structure of elements in the skiplist. That is, the highest level that an element appears at in a skiplist is equal to the number of successive “heads” seen when flipping a coin with some constant probabilty $p$ (usually $p=1/2$, but any probabilty $1/c$ where $c$ is a constant sufficies to achieve the asymptotic bounds). Given a skiplist with $n$ elements, the maximum height of any element is $O(\log n)$ in expectation and with high probability.
Operations. All operations in a skiplist start at the upper left sentinel (-$\infty$) at the highest level and traverse through the pointer structure in a left-to-right and top-to-bottom fashion. The skiplist is a self-balancing data structure and does not require pointer rotation to maintain its bounds. A skiplist with $n$ elements supports all point operations (insert, delete, find) in $O(\log n)$ time in expectation and with high probability.
For ease of understanding, given a node with key k at a given level $\ell$, let its pred (predecessor) element be the largest key strictly less than k, and let its succ (successor) element be the smallest key strictly greater than k at level $\ell$. Except for the beginning and end sentinels, all nodes in a skiplist have logical pred and succ elements which correspond to nodes in the skiplist.
To find a key k, the traversal searches left-to-right starting from the left sentinel on the highest level. Upon finding its succ node at that level, the search follows the down pointer of the pred node. The search continues in this way until the lowest level, where it scans left to right until it either finds k or does not find it and encounters some element greater than k.
The range operation is a direct extension of a find. Rather than terminating at the search key k or its succ element in the leaf level, the range search continues a left-to-right search through the leaf layer until length elements have been read, or the search reaches the rightmost sentinel.
Inserts are similar to finds in terms of traversal order, but must link in a new node containing some key k at each level that it is promoted to. Let $h\geq 0$ be the height that the newly inserted element is promoted to. One way to perform this structural modification is to keep track of the pred nodes of k at all levels during the downward traversal. After the search reaches the bottom level, the insert operation creates $h+1$ new nodes containing k and adjusts the next pointers at each level so that k->next = pred->next and then pred->next = k. Furthermore, the insert must update the down pointers in each internal node with the key k to the node containing k in the next lowest level.
2.2. B-skiplist structure and operations
Figure 3. Example of a B-skiplist with node size $B=4$.
Structure. The B-skiplist is a “blocked” version of a skiplist that contains multiple elements per node (Golovin, 2010). The elements are totally sorted at each level of the B-skiplist (by pointer structure and within nodes). Given a cache-line size $Z$ and desired node size $B=\Theta(Z)$, the B-skiplist’s promotion probability is generally set to $p=\Theta(1/B)$. In theory, the expected number of elements per node is $\Theta(B)$, but in a B-skiplist with $n$ elements, the maximum number of elements in a node is $\Theta(B\log n)$ w.h.p. (Bender et al., 2016, 2017). Golovin’s original theoretical paper on B-skiplists (Golovin, 2010) does not address how to handle this case in practice, but one straightforward solution is table doubling when the node becomes full.
The block structure depends on promoted height of each element in the B-skiplist. The header key of a node in a B-skiplist is the first (and smallest) element in that node. By construction, every header in a node at some level $\ell$ has been promoted to level $\ell+1$.
Figure 3 illustrates the node and pointer structure of a B-skiplist. Just as in a skiplist, each node has a next pointer pointing to the next node. Given a node x, all keys in x are smaller than the header key of x->next. Furthermore, each internal node (at level $\ell>0$) contains an array of $B$ down pointers (one per key in the node), each pointing to the corresponding node at the level below.
Bounds. In the external-memory model described in Section 1, given a cache-line size $Z$, a B-skiplist with $n$ elements and node size $B=\Theta(Z)$ supports finds and inserts in $\Theta(\log_{Z}(n))$ cache-line transfers in expectation (matching B-tree bounds). Furthermore, range queries with $r$ elements in the range take $\Theta(\log_{Z}(n)+r/Z)$ cache-line transfers in expectation. However, finds and inserts take $\Theta(\log n)$ cache-line transfers in the worst case w.h.p., matching the bounds of a non-blocked skiplist.
Operations. The left-to-right and top-to-bottom pointer traversal in a B-skiplist during finds and inserts is similar to the traversal in a skiplist. The main difference is that traversals in a B-skiplist must look at multiple elements within a single node and determine which down pointer to follow at an internal nodes.
To find a key k in a B-skiplist, the traversal begins at a curr node initialized to the upper left sentinel (with header -$\infty$), just as in the regular skiplist. The search then examines the header of curr->next. If it is less than k, the curr node is updated to curr->next. We repeat this left-to-right traversal until the header of curr->next is greater than k. At that point, we search within curr for the pred element and follow its down pointer. The search continues in this way until we reach the leaf level, at which point we determine if k is present. The main difference from a traditional skiplist is that the pred element may be in the same node as k.
To insert an element in a B-skiplist, we first perform a find-like traversal to find all the pred elements at each level. However, the node-modification operations during a B-skiplist insert are similar to those in a B-tree. Let $h$ denote the height at which the key to be inserted k is promoted to. If it not promoted (i.e., $h=0$), we can simply add it to the same node its pred element resides in and shift all subsequent elements in the node one slot down. However, if it is promoted (i.e., $h>0$), we must perform a split at levels $\ell=0,1,\ldots,h-1$. Let old_node be the node with pred. A node split in a B-skiplist creates a new_node with k as the header and copies all elements (and their down pointers, if the split occurs at an internal node) in old_node greater than pred (and also greater than k) after k. To link in the new node, we set new_node->next = old_node->next and old_node->next = new_node.
Open problem. Golovin’s paper on B-skiplists does not address the order of levels in which an element is inserted (i.e., starting from the top or the bottom) (Golovin, 2010). However, most traditional skiplist insertion algorithms insert elements in a “bottom-up” fashion - they link in new nodes starting from the leaf level up until the maximum level an element is to be promoted to (Pugh, 1990b, a; Herlihy et al., 2006, 2007). A similar algorithm for B-skiplists would perform a find to the bottom level, keeping track of the affected nodes along the way, and insert the element starting from the leaf level, performing splits and adjusting pointers on the way up as necessary to always maintain the inclusion invariant.
2.3. Concurrency control primitives
Hand-over-hand locking. Next, we will review the classical hand-over-hand (HOH) fine-grained locking scheme (also known as latch crabbing in databases) for sorted singly-linked lists (Herlihy et al., 2020), which we will be building upon in later sections. HOH traverses a list while holding at most two locks at a time, starting at the head node of the linked list, then acquiring the lock on the successor before releasing the lock on the predecessor. After locking a node curr, it is safe to access curr->next to either perform a comparison with a target key (e.g., for a search), or to see whether we have reached the end of the list. During inserts, a traversal must hold two locks at a time to link in a new node in the correct place between a node curr and curr->next. HOH locking can also be extended to skiplists because skiplists are simply towers of linked lists.
Reader-writer locks. Finally, we will review reader-writer locks (RW locks) (Courtois et al., 1971), as we will use them as the core synchronization primitive in the proposed concurrent B-skiplist. RW locks enable concurrent access for read-only operations but exclusive access for write operations. Given a RW lock, a thread can either call read_lock() to access it in shared read mode, or write_lock() to access it in exclusive write mode. Multiple threads can read the data in parallel if they all hold the lock in read mode, but all other threads must wait if a thread holds the lock in write mode.
RW locks are the foundation of many concurrency control protocols in B-trees, including the classical optimistic concurrency control (OCC) (Kung and Robinson, 1981).
3. Top-Down Insertion Algorithm
This section presents the top-down insertion algorithm that will form the basis of the concurrent B-skiplist in the next section. We will first describe the algorithm serially and show that it results in the same B-skiplist structure as the original “bottom-up” insertion algorithm. In the next section, we will introduce a corresponding top-down concurrency control scheme. Furthermore, we will show that the proposed insertion algorithm achieves the same asymptotic runtime bounds of the theoretical B-skiplist. For simplicity, we will describe the algorithm for keys only, but storing associated values involves only updating the leaf level. We will initially describe the insertion algorithm using logical nodes for simplicity and then explain how to adapt it with fixed-size physical nodes.
Figure 4. Example of a top-down insertion of key k = 30 and height $h=1$ into a B-skiplist with node size $B=4$. The dotted box represents the current node visited during the traversal, the blue dashed lines denote the pointers followed during the top-down traversal, and the blue cells represent the cells written during the insertion.
Description. At a high level, the main change in the proposed “top-down” insertion algorithm compared to existing “bottom-up” algorithms is the order in which new elements are added to the skiplist. Although the distinction about the order of insertion into levels is subtle, it has important implications for the traversal direction, and as we shall see in Section 4, for CC mechanisms.
The goal of the top-down algorithm is to complete insertions in one pass of the skiplist without revisiting nodes by taking advantage of the skiplist property that the height at which an element is promoted to is determined upfront and is independent of the current structure. In contrast to promotions in a B-tree, which depend on the current structure and fullness of the nodes, element promotions in a B-skiplist depend only on a sequence of random coin flips.
Rather than traversing down to the leaf level and then linking in elements from bottom to top, as in most skiplist insertion algorithms, we propose to insert elements starting from the highest level to which they are promoted and ending with the leaf level. Therefore, an insertion is finished once the traversal reaches the leaf level and performs a write at the relevant node.
Suppose we are inserting an element k that will be promoted to level $h\geq 0$ as determined by random coin flips. We traverse the B-skiplist from top to bottom starting at the leftmost sentinel at the top level as described in Section 2. If the traversal is currently on some level $\ell>h$, the element k will not appear on $\ell$, so the search order is exactly the same as an original B-skiplist find. By following next and down pointers, we will eventually reach level $h$, where we will write k into the same node as its predecessor pred at level $h$. If $h=0$, the insertion is complete.
The main challenge comes when the element k is promoted to level $h>0$ because the corresponding down pointer should point to a new node (due to a split) at level h1h-1 that has not yet been created in a naive top-down traversal. Recall from Section 2 that in a B-skiplist, every header element in a node at level $i$ must have been promoted to level $i+1$. It is clear which node to set the down pointer to with a “bottom-up” algorithm, because the corresponding node has been created at level h1h-1 before the down pointer at level $h$ needs to be set. However, with the proposed “top-down” algorithm, the node with the new element k has not yet been created at level h1h-1 since we are going from $h$ down to 0.
We can resolve the issue by again exploiting the property of B-skiplist insertions that the height of every key is determined upfront to allocate all the nodes in advance that will be created during an insertion. That is, when an element is promoted to height $h$, we can allocate $h$ new nodes with k as the header at the start of any insertion before any interaction with the skiplist at all. If $h>1$, we can link these preallocated nodes together in a stack via down pointers in non-leaf nodes. Therefore, we can fill in the appropriate slot in the down pointer array at level $h$ with a pointer to the top of the preallocated stack of new nodes.
To determine where to splice in the other new nodes in levels 0 to h1h-1, we continue the traversal level-by-level. Let us consider the levels in turn, starting with level h1h-1. Just as in the higher levels, the traversal will follow the down pointer corresponding to the prev element in level $h$ which will point to some node at level h1h-1. At level h1h-1, the traversal will proceed left-to-right until we find some node x such that x->header < k but x->next->header > k. Let n be the preallocated node destined to be at level h1h-1. Elements greater than k in node x should be moved to n. It is then linked into the B-skiplist by setting n->next = x->next and x->next = n. The insertion then proceeds in this way until it reaches the leaf level by following the down pointer of the last element in x, which is the largest element at that level less than k. The element k will appear in the correct nodes with the correct structure at the end of this top-down method with splits on the way down.
Figure 4 contains a worked out example of a top-down insertion. Notice how in the example, since the element is promoted to height $h=1$, the corresponding node at level 0 is allocated upfront and linked in with down and next pointers at levels 1 and 0, respectively.
Correctness and bounds. The remainder of the section will describe the differences resulting from the top-down insertion algorithm and how they do not affect the B-skiplist’s correctness or asymptotic time bounds.
The main change with the proposed top-down insertion algorithm is a slight relaxation of the inclusion invariant, or the property that every element present in level $h$ must also be present at all lower levels, during insertions. It still preserves the property upon every insertion’s completion by linking in the preallocated new nodes via down and next pointers.
The inclusion invariant is necessary for the B-skiplist to achieve its asymptotic bounds, because the bounds come from the expected number of levels and the expected number of nodes traversed in each level. These properties are derived from the probabilistic coin flips that determine each element’s height in the B-skiplist. However, after each insertion, the pointer structure in the B-skiplist with the top-down insert algorithm is identical to what it would have been with a bottom-up algorithm that links in nodes starting from the leaf level. Therefore, both the number of levels and the number of nodes traversed per level are unaffected.
Fixed-size nodes. As mentioned in Section 1, due to the B-skiplist’s randomized structure, any practical B-skiplist implementation must handle the case where the number of elements in a node exceeds some fixed-size array allocation. The number of keys that are supposed to be in a node may exceed a fixed bound, depending on the sequence of coin flips. Given a B-skiplist with $n$ elements and promotion probability $p=1/B$, the worst-case number of elements in a given node is $\Theta(B\log n)$ w.h.p. Therefore, it is highly likely that there will be nodes that exceed any fixed size $\Theta(B)$.
The naive solution of table doubling in nodes enables all keys meant for a node to fit in the corresponding array, but leads to suboptimal performance of insertions. As mentioned in Sections 1 and 2, w.h.p., there is a node in the B-skiplist with $k=\Theta(B\log n)$ elements, so the worst-case insertion cost in a B-skiplist is $\Theta(\log n)$ cache-line accesses w.h.p.
To alleviate this issue, we modify the B-skiplist design to require fixed-size node allocations, which may potentially result in node overflow splits (i.e., node splits due to overflow) in addition to promotion splits due to randomization. If nodes were allowed to grow arbitrarily large, the worst-case number of elements moved (i.e., shifted to maintain sorted order) during an insert is $\Theta(B\log n)$, which would take $\Theta(\log n)$ cache misses. However, by requiring that the nodes have at most $B$ elements, the number of element moves in any node is at most $B$.
The nodes created from overflow splits do not affect the correctness of inserts or searches, as operations in the B-skiplist still follow a left-to-right and top-to-bottom traversal order. As mentioned in Section 1, the choice of fixed-size nodes does not affect the theoretical bounds, because a query would still have to perform $\Theta(\log n)$ cache-line transfers in the worst case w.h.p. However, minimizing the cost of inserts is important for practical efficiency, as element moves must be linear in the node size, while queries can skip over parts of the node e.g., via binary search.
4. Top-down concurrency control
This section presents a single-pass top-down concurrency control scheme for B-skiplists based on reader-writer locks and the top-down insert scheme described in Section 3. The goal of the proposed scheme is simplicity in both the number of top-down traversals and the number of locks held at a time. Specifically, we will show that each operation only needs to make a single root-to-leaf traversal. Furthermore, this traversal only holds a constant number of locks in at most two levels of the B-skiplist at a time.
Concurrent finds and range queries. Let us start with how to implement concurrent finds and range queries with RW concurrency as an intermediate step to understanding concurrent inserts in B-skiplists. Since finds and range queries are read-only operations, they only need to acquire locks in reader mode [^5] Reader locks are necessary in mixed insert-query workloads.. Queries begin on the highest level at the left sentinel and proceed in a hand-over-hand fashion left-to-right, as described in the HOH scheme for linked lists in Section 2. When the search reaches the node with the appropriate prev element, it acquires the child node at the next lowest level via the down pointer using HOH locking in a top-down fashion. Searches proceed left-to-right within a level and top-to-down to move between one level at a time until the query reaches the appropriate node in the leaf level that should contain the target key. For point finds, the search is then complete and can release all locks. In contrast, range queries acquire locks left-to-right at the leaf level in a HOH fashion until the range is exhausted or the search reaches the end of the skiplist.
To recap, both concurrent finds and range queries acquire RW locks in read mode in a left-to-right and top-to-bottom order in HOH fashion. That is, a thread holds at most two locks at a time.
Figure 5. Example of the insertion of element with key k = 30 and height $h=1$ into a B-skiplist with node size $B=4$ using the top-down CC scheme. The locks with R and W denote acquiring the lock in read and write mode, respectively.
Concurrent inserts. Next, we will introduce the proposed top-down concurrency control scheme for inserts. At a high level, insertions follow a similar left-to-right and top-to-bottom traversal order to queries. However, inserts raise additional challenges, since they potentially require structural modification operations (i.e., splits) if elements are promoted to higher levels.  Figure 5 illustrates a worked example of the sequence of reader-writer lock acquisitions and node updates with the top-down concurrency control protocol.
Just as in the insertion algorithm from Section 3, the proposed top-down concurrency control scheme leverages the randomized property of skiplists to complete insertions in one pass through the data structure and to acquire writer locks only at the levels where writes will occur. That is, it relies on the observation that the level to which an element is promoted to in a B-skiplist depends only on a sequence of random coin flips and importantly, is can be determined upfront independently of the current structure of the skiplist.
Suppose that we are inserting some key k that will be promoted to level $h$. Furthermore, suppose that we have preallocated any new nodes that will be spliced into the skiplist as described in Section 3. Since they are not currently linked in the skiplist, taking their write locks will not delay any other threads.
Inserts begin at the highest level and proceed just like queries for all levels greater than $h$. Since there will be no writes until level $h$, the traversal only needs to acquire locks in read mode in the left-to-right and top-to-bottom traversal until it reaches level $h$.
Once it has reached the appropriate node with the correct down pointer in level $h+1$, the insert then acquires the lock on the child node in write mode on level $h$. At level $h$, the traversal will continue left-to-right in a HOH fashion taking locks in write mode until we reach the node where k should be inserted (i.e., some node x such that x->header < k and x->next->header > k). Since we have the write lock on x, we can directly insert k in the appropriate slot. If $h=0$, the insert is completed, and we can release all of the locks.
If there need to be splits at lower levels (i.e., $h>0$), we use HOH locking from top-to-bottom and left-to-right. Consider the case of a split at level h1h-1. To find the starting point of the traversal at level h1h-1, we first take the write lock on the node pointed to the by down pointer associated with the prev element in level $h$. Just as in the single-threaded case, the traversal will then proceed left-to-right until we find the two nodes that the new preallocated node will need to be inserted between. However, in the concurrent case, we take writer locks in HOH fashion left-to-right for thread safety. The actual split mechanism is unchanged from the serial case, and this process is repeated until the leaf level.
Integration with fixed-size nodes. So far, we have described the CC mechanism on logical nodes, but overflow splits do not affect the guarantees of HOH locking. An insert still only needs to acquire a constant number of locks at any time. The CC protocol in the case of overflow splits is even simpler than the case of promotion splits. Recall that if an element is promoted, the top-down scheme acquires at most three locks at once to perform splits: one to update the down pointer at some level $\ell$ and two to splice in the new node between two existing nodes at level $\ell-1$). In overflow splits, there is no need for the lock to be held on level $\ell$ because there is no down pointer to the new node, so we only need to hold the two locks on the nodes that we are splicing the new node between.
Correctness and deadlock-freedom. The correctness of the proposed HOH-based locking scheme follows directly from existing theory about HOH locking in linked lists and skiplists (Herlihy et al., 2020). To insert an element into a node without a split, just acquiring the write lock on that node is sufficient. To split a node and insert a new node between two existing ones, acquiring the write lock on both the predecessor node and its next node is sufficient. There is a slight relaxation of the maximum number of locks held at once from two to three during splits. However, locks are held on at most two levels at once.
Finally, the proposed top-down concurrency scheme is deadlock-free because there is a total ordering on locks from left-to-right within levels and then top-to-bottom between the levels, thereby avoiding circular wait.
5. Evaluation
This section evaluates the proposed concurrent B-skiplist on the YCSB (Cooper et al., 2010) compared to several state-of-the-art concurrent skiplist- and tree-based indices. As mentioned in Section 1, we evaluate all indices in terms of both throughput and latency.
Result summary. At a high level, the B-skiplist achieves $2\times$$9\times$ higher throughput and $3.5\times$$103\times$ lower latency than non-cache-optimized concurrent skiplists. Furthermore, the B-skiplist achieves $0.9\times$$1.7\times$ throughput on point workloads and $0.7\times$$7.5\times$ throughput on range workloads compared to tree-based indices.
Systems setup. All experiments were run on a server with 64-core 2-way hyperthreaded Intel Xeon Gold 6338 CPU @ 2.00GHz with 1008 GB of memory. The server has a 3 MiB L1 data cache, a 2 MiB L1 instruction cache, a 80 MiB L2 cache and a 96 MiB L3 cache. We ran all experiments with $64$ physical cores and $128$ hyperthreads.
All times are the median of 5 trials after one warm-up trial. To measure latency, each thread measures the average time taken for a batch of ten operations [^6] We measure the average of 10 operations instead of individual operation to preserve the contention between threads. and stores it in a thread-safe vector. This allows us to sort and calculate the latency at each percentile after running each benchmark.
We measured the cache misses in Table 2 with perf.
Workloads.  Table 2 presents details of the core workloads from the YCSB (YCSB, [n. d.]). We tested workloads [^7] We omit workload D from YCSB because it benchmarks the read-latest operation, which is not the focus of this work. A, B, C, and E from the core YCSB workloads generated with RECIPE (Lee et al., 2019).
We follow the standard YCSB procedure which consists of two phases: 1) the load phase, where elements are inserted into the data structure, and 2) the run phase, where operations are executed according to the workload’s find/insert ratio. Each workload consists of 100 million (100M) elements inserted during the load phase, followed by 100M operations executed during the run phase. All operations in each phase are performed concurrently, but the phases are performed one after another.
We evaluate each workload under both uniform random and zipfian distributions in the run phase. In the uniform workload, the elements in both the load and run phases are generated from a uniform distribution. In the zipfian workload, the elements in the load phase are generated from a uniform distribution, while the elements in the run phase are generated from a zipfian distribution. We omit the results on the zipfian distirbution due to space limitations, but the results were similar (on average within $20\%$) to the uniform distribution.
Workload Description
Load 100% inserts from empty
A 50% finds, 50% inserts
B 95% finds, 5% inserts
C 100% finds
E 95% short range iterations (max_len = 100), 5% inserts
Table 2. YCSB workload descriptions.
B-skiplist setup. We implemented the concurrent B-skiplist in C++ with an open-source reader-writer lock library (par, [n. d.]). The test driver executes concurrent operations using pthreads (Nichols et al., 1996). We ran the B-skiplist with 8-byte keys and 8-byte values (for 16-byte key-value pairs). We set the max height of the B-skiplist to 5 in our tests.
We compiled the B-skiplist using g++ 11.4.0 with -O3.
B-skiplist sensitivity analysis. We performed a parameter sweep over promotion probability and node size in the B-skiplist to empirically determine which settings yield the best performance on the YCSB workloads. Golovin’s theoretical paper on B-skiplists proposes a scaling factor, some constant $c$, on the promotion probability for $p=1/cB$. Therefore, for each tested node size we also experiment with $c\in\{0.5,1.0,2.0\}$. In theory, any constant $c$ should suffice to achieve the theoretical randomized B-skiplist bounds (Golovin, 2010).
Based on the results of this sensitivity analysis, we set the node size in the B-skiplist to 2048 bytes (i.e., 128 key-value pairs) and $c=0.5$, for promotion probability $p=1/(0.5\times 128)=1/64$.
5.1. Comparison to skiplist-based indices
In this section, we evaluate the B-skiplist against No Hot Spot Skip List  (Crain et al., 2013), Java ConcurrentSkipListMap (Java, [n. d.]), and Facebook Folly’s ConcurrentSkipList  (fol, [n. d.]) on the YCSB workloads (YCSB, [n. d.]) and report the results in Figures 1 and 6.
Systems setup. Java ConcurrentSkipListMap (JSL) implements a concurrent skiplist using a tree-like 2D linked skiplist (Michael, 2002).All operations except range queries are done natively. We implement range queries with the subMap interface.
Facebook’s Folly (FLY) library provides a C++ concurrent skiplist [^8] https://github.com/facebook/folly/blob/main/folly. We used the native interface for all point operations and the iterator interface for range queries (since it does not have native support). Folly’s skiplist does not support values so we store only the keys.
No Hot Spot Skip List999 https://github.com/wangziqi2016/index-microbench/tree/master/nohotspot-skiplist (NHS) is a concurrent, lock free skiplist in C++ that relies on a background adaptation thread to maintains the structure of the skiplist and manage garbage collection. This includes rebalancing the upper index level to ensure traversals can be done in $O(\log n)$ time. NHS takes in a sleep time parameter that determines how frequent the background thread checks the index and modifies it. In the load phase, we set this parameter to a relatively small value (100 microseconds) to ensure the index is frequently balanced. After the load phase, we must wait for the background thread to balance the height of the tree to $\lg n$ to ensure that operations in the run phase achieve the desired performance. In the run phase, we set the sleep time higher to 1 second to ensure the background thread does not stall operations. We do not count the rebalance time between the load and run phases.
We compiled FLY and NHS using g++ 11.4.0 with -O3 and Java ConcurrentSkipListMap using javac 11.0.25.
[…]
Figure 6. Latencies of skiplist-based indices at different percentiles in YCSB workload A with uniform random keys.
Discussion.  Figures 1 and 6 illustrate the throughputs and various percentile latencies of the skiplists.
[…]
Additionally, as shown in Figure 6, the B-skiplist achieves lower latency across all workloads in all tested percentiles (50%, 90%, 99%, and 99.9%). The most competitive non-blocked skiplist is the one from Folly, which has at least $3.5\times$ and often at least $3.7\times$ higher latency compared to the B-skiplist in the different percentiles.
[…]
5.2. Comparison to tree-based indices
Load A B C E 00.50.5111.51.511Workload Normalized performance B-skiplistB-treeMassTree
Figure 7. Normalized throughput (ops/s) of tree-based indices relative to the B-skiplist with uniform random keys.
Systems setup. We compare the B-skiplist with a high-performance concurrent B+-tree 101010 https://github.com/wheatman/BP-Tree/tree/main/tlx-plain/container (Xu et al., 2023) (a common B-tree variant) based on optimistic concurrency control (Kung and Robinson, 1981). The default configuration for the B+-tree sets $\texttt{node\_size}=1024$ bytes. Both the concurrent B-skiplist and concurrent B+-tree use the same RW lock.
We also compare the B-skiplist with Masstree111111 https://github.com/kohler/masstree-beta (Mao et al., 2012), a cache friendly B+-tree variant. It utilizes an optimistic concurrency scheme as well. There is native support for all YCSB operations.
We compiled the B+-tree and Masstree with g++ 11.4.0 and -O3.
Discussion.  Figures 7 and 8 compare the B-skiplist to tree-based indices on throughput and latency, respectively. We tested both uniform and zipfian datasets but only illustrate the performance on uniform as the results are similar.
[…]
Figure 8. Percentile latencies of the B-skiplist and tree-based indices on YCSB workload A with uniform random keys.
On the other hand, the B-tree achieves about $1.4\times$ higher throughput than the B-skiplist on range queries (Workload E). As mentioned earlier, although the B-skiplist has $\Theta(B)$ elements per node in expectation, the number of elements actually in a logical node can vary by up to a factor of $O(\log n)$. Therefore, the average density (number of elements in a node) is lower in the B-skiplist compared to the B+-tree, which deterministically splits nodes when they become full. To measure this difference, we counted the number of nodes at the bottom level traversed in both the B-skiplist and B+-tree during workload E. On average, the B-skiplist accesses about $2$ nodes per range query while the B+-tree accesses only about $1.5$ nodes per range query. We note that in-memory indexes are traditionally optimized for point (OLTP) workloads and range queries are often an optional function. Future work involves improving range queries in B-skiplists by improving the average node density.
5.3. Strong Scaling
Figures 9 and 10 shows the scaling performance of B-skiplist, Folly, B-tree, JSL, NHS, and Masstree on YCSB workload A and C with uniform random keys. NHS scaling starts from 2 threads since it requires at least two threads to run.
[…]
124816326412810102020303040405050Num. ThreadsSpeedupB-skiplistFollyJSLB-treeMasstreeNoHotSpot
Figure 9. Strong scaling of various systems in terms of throughput on YCSB workload A.
12481632641281010202030304040505060607070Num. ThreadsSpeedupB-skiplistFollyJSLB-treeMasstreeNoHotSpot
Figure 10. Strong scaling of various systems in terms of throughput on YCSB workload C.
6. Conclusion
We present the B-skiplist, a high-performance concurrent in-memory index based on the skiplist data structure. The proposed concurrent B-skiplist adapts the theoretical description of a B-skiplist (Golovin, 2010) for practical considerations to minimize data movement and mitigate the probabilistic worst case of element moves. To take advantage of parallel resources, we propose a top-down insertion algorithm that completes insertion in one pass and a corresponding simple yet effective CC scheme. The B-skiplist inherits the simple structure that makes the skiplist a popular choice for in-memory indexing.
The empirical evaluation demonstrates that the B-skiplist achieves between $2\times$$9\times$ higher throughput and between $3.5\times$$103\times$ lower 99th percentile latency compared to popular state-of-the-art concurrent skiplist implementations such as those from Facebook’s folly library and the Java concurrent skiplist library. These results suggest that the B-skiplist is a good candidate for high-performance in-memory indexing because it resolves locality issues in skiplists while minimizing CC overhead.
For future work, we plan to integrate the B-skiplist into key-value stores like RocksDB and LevelDB to evaluate its impact on application performance. Additionally, its high cache locality makes it well-suited for disk-based indexes. With its strong theoretical guarantees and practical efficiency, we anticipate that B-skiplist could match or even surpass traditional skiplists in modern databases.
Acknowledgments
This research is funded in part by NSF grant OAC 2339521 and 2517201.
Appendix A Pseudocode
Algorithm 1 Insertion Algorithm
1:SkipList SLSL and a pair key $K$ value $V$
2:$K$,$V$ inserted into SLSL
[…]
4:current $\leftarrow$ starting sentinel
5:current.lock(h, MAXHEIGHT)
6:for level $\leftarrow$ MAXHEIGHT to 0 do
[…]
8:  while current.next_header <=<= key do
9:   current.next.lock(h, level)
10:   previous $\leftarrow$ current
11:   current $\leftarrow$ current.next
12:   previous.unlock
13:  end while
14:  rank, found $\leftarrow$ current.find_key($K$)
15:  if found then
16:   if map then
17:     current.write(K,V) at rank
18:   else if set then
19:     current.unlock and then Exit
20:   end if
21:  else
22:   if h ==== level then
23:     if current is overflowing then
24:      new_node and then new_node.lock(h, level)
25:      new_node.next $\leftarrow$ current.next
26:      current.next $\leftarrow$ new_node
27:      half $\leftarrow$ current.num_elements/2
28:      current.split(new_node, half)
29:      if rank +1<=+1<= current.num_elements then
30:       new_node.unlock
31:      else
32:       current.unlock
33:       current $\leftarrow$ new_node
34:       rank $\leftarrow$ rank - current.num_elements
35:      end if
36:     end if
37:     current.insert(K,V) at rank++1
38:     if level >>then
39:      current.insert_child(rank++1)
40:     end if
41:   else if h ¡ level then
42:     new_node
43:     new_node.next $\leftarrow$ current.next
44:     current.next $\leftarrow$ new_node
45:     new_node.insert(K,V) at 0
46:     current.split(new_node, rank++1)
47:     if level >>then
48:      current.insert_child(rank++1)
49:     end if
50:   end if
51:  end if
52:  if level >>then
53:   previous $\leftarrow$ current
54:   current.child(rank).lock(h, level)
55:   current $\leftarrow$ current.child(rank)
56:   previous.unlock
57:  end if
58:end for
Algorithm 2 Lock Logic
1:Promotion level $l$, current traversal level $h$, current node $n$
2:Either read or write locks the current node
3:if l ¡ h then
4:  n.read_lock
5:else
6:  n.write_lock
7:end if
Appendix B Deadlock Freedom
Deadlock-freedom. We will show that the proposed concurrent B-skiplist is deadlock-free. The most popular way to avoid deadlock is to ensure that locks are acquired in a total order, thereby avoiding circular wait. In the proposed top-down concurrency control protocol, there is a total ordering on the locks from left-to-right within a level, and then from top-to-bottom in the levels. That is, the lowest ordered node is the left -$\infty$ sentinel at the highest level, and the highest ordered node is the right $\infty$ sentinel at the leaf level. The left -$\infty$ sentinel at some level $\ell$ is ordered directly after the right $\infty$ sentinel at level $\ell+1$. The insert and find algorithms acquire locks according to this total ordering in a left-to-right and top-to-bottom fashion, thereby avoiding deadlock.
Appendix C Zipfian Key Figures
Load A B C E 11551010Workload Normalized throughput NHSFolly SLJava SLB-skiplist
Figure 11. Normalized throughput (ops/s) of skiplist-based indices relative to the No Hot Spot Skip List(NHS)  (Crain et al., 2013) under Zipfian key distribution.
Load A B C E 00.50.5111.51.511Workload Normalized performance B-skiplistB-treeMassTree
Figure 12. Normalized throughput (ops/s) of tree-based indices relative to the B-skiplist with zipfian keys.
[…]
Figure 13. Latencies of tested indices at different percentiles in YCSB workload A with zipfian keys.
Appendix D Data Tables
100% finds 100% inserts
Bytes Elts $c$ TP DFB 90% DFB 99% DFB 99.9% DFB TP DFB 90% DFB 99% DFB 99.9% DFB
0.5 28.7 0.6 2.41 0.5 2.8 0.5 4.55 0.4 13.2 0.4 4.09 0.5 15.45 0.2 284.57 0.0
512 32 1 42.5 0.8 1.44 0.8 1.65 0.8 2.16 0.9 21.2 0.7 3.1 0.7 4.22 0.7 7.09 0.6
2 37.7 0.7 1.66 0.7 1.94 0.7 2.44 0.8 14.8 0.5 4.95 0.4 11.23 0.3 20.97 0.2
0.5 49.8 1.0 1.23 1.0 1.45 0.9 2 1.0 29.2 1.0 2.12 1.0 3.09 0.9 4.69 1.0
1024 64 1 49.0 1.0 1.28 0.9 1.52 0.9 2.12 0.9 22.2 0.7 2.62 0.8 5.21 0.6 10.81 0.4
2 42.1 0.8 1.57 0.7 1.87 0.7 2.55 0.8 14.9 0.5 4.96 0.4 7.74 0.4 12.58 0.4
0.5 50.6 1.0 1.17 1.0 1.37 1.0 1.92 1.0 30.2 1.0 2.02 1.0 2.91 1.0 5.3 0.8
2048 128 1 46.6 0.9 1.31 0.9 1.56 0.9 2.2 0.9 22.9 0.8 2.72 0.7 3.91 0.7 6.39 0.7
2 43.8 0.9 1.5 0.8 1.76 0.8 2.48 0.8 14.1 0.5 5.35 0.4 9.88 0.3 15.81 0.3
0.5 45.1 0.9 1.34 0.9 1.57 0.9 2.27 0.8 25.7 0.9 2.46 0.8 3.32 0.9 4.48 1.0
4096 256 1 43.2 0.9 1.44 0.8 1.69 0.8 2.64 0.7 20.5 0.7 3.25 0.6 5.1 0.6 7.94 0.6
2 36.3 0.7 1.79 0.7 2.17 0.6 3.48 0.6 9.6 0.3 10.52 0.2 21.6 0.1 34.06 0.1
0.5 37.7 0.7 1.67 0.7 2.02 0.7 3 0.6 17.8 0.6 3.81 0.5 5.14 0.6 6.62 0.7
8192 512 1 31.8 0.6 2.08 0.6 2.54 0.5 4.32 0.4 13.8 0.5 5.76 0.4 10.95 0.3 16.81 0.3
2 33.2 0.7 1.98 0.6 2.35 0.6 3.53 0.5 10.9 0.4 7.78 0.3 13.86 0.2 19.71 0.2
[…]
Uniform Zipfian
Metric YCSB BSL JSL JSL/BSL NHS NHS/BSL FLY FLY/BSL BSL JSL JSL/BSL NHS NHS/BSL FLY FLY/BSL
Load 20.6 10.0 0.5 1.9 0.1 9.7 0.5 20.3 9.8 0.5 1.9 0.1 9.7 0.5
A 42.8 9.2 0.2 8.6 0.2 12.8 0.3 48.7 10.3 0.2 13.5 0.3 13.1 0.3
TP B 75.3 11.5 0.2 8.5 0.1 24.5 0.3 89.0 13.8 0.2 11.3 0.1 28.7 0.3
C 74.1 11.7 0.2 10.5 0.1 23.0 0.3 83.3 15.6 0.2 21.3 0.3 29.3 0.4
E 45.3 5.5 0.1 6.1 0.1 5.7 0.1 54.9 6.2 0.1 8.9 0.2 5.5 0.1
Load 1.5 7.2 4.7 52.4 33.8 10.8 7.0 1.5 7.6 4.9 52.4 33.8 10.8 7.0
A 1.4 10.2 7.3 3.5 2.5 7.4 5.3 1.3 9.8 7.7 3.3 2.6 8.8 6.9
50% B 1.2 9.6 8.2 3.4 2.9 4.6 3.9 1.0 9.5 9.5 2.8 2.8 3.9 3.9
C 1.2 9.8 8.5 3.5 3.0 4.3 3.8 1.0 8.4 8.7 2.7 2.8 3.7 3.8
E 1.8 13.8 7.6 12.3 6.7 16.7 9.2 1.6 13.5 8.6 9.5 6.0 20.3 12.9
Load 1.9 9.8 5.2 112.1 59.1 13.8 7.3 1.9 10.0 5.3 112.1 59.1 13.8 7.3
A 1.6 13.4 8.2 4.3 2.6 4.3 2.6 1.5 12.9 8.3 3.9 2.5 3.9 2.5
90% B 1.3 12.6 9.6 4.5 3.4 5.2 3.9 1.2 12.0 10.2 3.7 3.2 4.6 3.9
C 1.3 12.6 9.6 4.7 3.6 4.9 3.8 1.1 11.1 9.8 3.6 3.1 4.3 3.8
E 2.1 20.5 9.9 20.0 9.7 37.2 17.9 1.8 18.5 10.1 16.2 8.8 27.8 15.1
Load 2.5 9.8 3.9 257.5 103.2 16.9 6.8 2.5 10.0 4.0 257.5 103.2 16.9 6.8
A 2.2 15.9 7.3 91.8 42.3 10.8 5.0 2.1 15.3 7.4 44.4 21.6 13.5 6.6
99% B 1.7 14.4 8.4 154.2 89.9 6.0 3.5 1.5 13.8 9.2 77.5 51.6 5.5 3.6
C 1.5 14.5 9.4 150.0 97.5 5.6 3.7 1.3 13.0 9.7 76.1 57.0 5.1 3.8
E 2.5 29.9 12.1 165.5 67.2 47.2 19.2 2.2 22.7 10.3 88.5 40.0 33.7 15.3
Load 51.3 35.0 0.7 982.4 19.1 20.0 0.4 51.3 32.6 0.6 982.4 19.1 20.0 0.4
A 3.2 26.0 8.1 91.8 28.5 12.7 3.9 3.0 51.4 17.3 44.4 14.9 15.6 5.2
99.9% B 2.3 26.0 11.1 154.2 65.9 9.6 4.1 2.1 51.4 24.1 77.5 36.4 9.2 4.3
C 2.2 56.8 26.0 150.0 68.8 9.2 4.2 2.0 16.2 8.3 76.1 38.8 8.6 4.4
E 2.2 536.8 246.1 150.0 68.8 55.9 25.6 2.0 330.0 168.3 76.1 38.8 41.3 21.1
[…]
Uniform Zipfian
OBT MT OBT MT
Metric YCSB OBT BSL MT BSL OBT BSL MT BSL
Load 15.5 0.8 9.6 0.5 14.5 0.7 9.6 0.5
A 31.2 0.7 24.0 0.6 30.5 0.6 21.7 0.4
TP B 75.3 1.0 60.7 0.8 90.7 1.0 72.9 0.8
C 81.6 1.1 69.7 0.9 88.7 1.1 94.8 1.1
E 64.0 1.4 6.0 0.1 68.9 1.3 6.9 0.1
Load 1.9 1.3 3.6 2.4 19.5 12.6 36.4 23.5
A 1.4 1.0 2.6 1.8 12.5 9.7 27.1 21.2
50% B 1.1 0.9 1.5 1.3 8.7 8.7 12.5 12.5
C 1.0 0.9 1.5 1.3 8.4 8.7 12.1 12.6
E 1.3 0.7 21.4 11.8 10.7 6.8 181.7 116.0
Load 12.6 6.6 5.6 3.0 125.6 66.2 56.2 29.6
A 2.0 1.2 3.3 2.0 17.0 11.0 34.9 22.6
90% B 1.2 0.9 1.6 1.2 10.3 8.8 15.0 12.8
C 1.1 0.9 1.6 1.3 9.8 8.7 14.1 12.4
E 1.4 0.7 26.6 12.8 12.6 6.9 232.3 126.5
Load 33.2 13.3 160.4 64.3 331.7 132.9 1604.5 643.1
A 13.3 6.1 77.0 35.5 122.9 59.8 857.1 417.3
99% B 3.2 1.9 2.4 1.4 18.2 12.1 22.2 14.8
C 1.3 0.9 2.2 1.5 11.4 8.5 20.0 15.0
E 2.7 1.1 30.7 12.5 30.3 13.7 274.5 124.2
Load 122.6 2.4 410.4 8.0 1226.4 23.9 4104.3 79.9
A 24.7 7.7 187.6 58.2 256.5 86.2 1971.9 662.8
99.9% B 12.0 5.1 3.0 1.3 97.8 45.9 24.2 11.3
C 2.0 0.9 2.4 1.1 17.0 8.6 23.5 12.0
E 12.2 5.6 35.2 16.2 109.8 56.0 307.1 156.6
[…]
Table 3 contains the results of the parameter sweep that was used to determine the optimal parameter. The experiment was ran on 64 threads in a single NUMA socket.
Table 4 and  Table 5 contains the data used to generate figures 1,  6,  7,  8,  11,  12, and  13.
Appendix E Artifact Instructions
The B-skiplist implementation is available at https://github.com/Ratbuyer/bskip_artifact, including all the tested systems in Section 5. The README files in the repository contain compiling and running instructions.